• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Michaels

    Log entry time 07:51:16 on April20,2007

    Entry number 199701

    keyword=owl shift summary

    Doing kinematics #9 for Qsq = 50 fm-2
    I verified the magnet settings and angles, and that the
    trigger was downloaded by last shift.

    Runs 2682 - 2690 production on He4

    04:10 Switch to dummy target. Reduce current to 40 uA.

    Run 2691 with 20 cm Al dummy

    05:30 Switch back to He4. Go to 82 uA (all we can get)

    Run 2692 - 2696 production on He4

    Run 2696 has 800K events. The beam was taken away at 07:50
    for shutdown -> open house.